# Leveraging Symmetry Merging in Pauli Propagation

### v1

@@ B Symmetry and Group Theory
In this self-contained section, we review the terminologies that we use from group theory, which can be found in standard texts [armstrong1997groups, rotman2012introduction].
@@
Lemma 1 (Burnside’s lemma [armstrong1997groups]).
@@
Corollary 1 (Pauli representatives cardinality [armstrong1997groups]).
@@
This is also called the tetrahedral number $\textrm{Te}_{n}=\frac{n(n+1)(n+2)}{6}$ [schatzki2024theoretical]. For large $n$, the number of orbit representatives scales as $\Theta(n^{3})$, leading to $r\in\Theta\left(\frac{n^{3}}{4^{n}}\right)$ improvement in terms of space complexity.
@@ D Computational Complexity
Finding the largest orbit for an arbitrary group is an open question, to the best of our knowledge [qian2021large], hence this lower bound is not necessarily achievable. Nevertheless, for most common symmetry groups appearing in the context of actions on Pauli strings in quantum simulations, our lower bound in Observable 2 is usually satisfied. In the rest of the section, we analyze the time complexity of our algorithm and comment on its relations to practical simulations.
@@ E.2 Practical subroutines: lowest integer mapping
where multiplication of Pauli operators corresponds to addition modulo two in the symplectic vector space [rudolph2025pauli]. Here, we follow the base-$4$ numeral conventions as Ref. [rudolph2025pauli], to which we refer the reader for further implementation details.

### v2

@@ B Symmetry and Group Theory
In this self-contained section, we review the terminologies that we use from group theory, which can be found in standard texts [7, 51].
@@
Lemma 1 (Burnside’s lemma [7]).
@@
Corollary 1 (Pauli representatives cardinality [7]).
@@
This is also called the tetrahedral number $\textrm{Te}_{n}=\frac{n(n+1)(n+2)}{6}$ [55]. For large $n$, the number of orbit representatives scales as $\Theta(n^{3})$, leading to $r\in\Theta\left(\frac{n^{3}}{4^{n}}\right)$ improvement in terms of space complexity.
@@ D Computational Complexity
Finding the largest orbit for an arbitrary group is an open question, to the best of our knowledge [46], hence this lower bound is not necessarily achievable. Nevertheless, for most common symmetry groups appearing in the context of actions on Pauli strings in quantum simulations, our lower bound in Observable 2 is usually satisfied. In the rest of the section, we analyze the time complexity of our algorithm and comment on its relations to practical simulations.
@@ E.2 Practical subroutines: lowest integer mapping
where multiplication of Pauli operators corresponds to addition modulo two in the symplectic vector space [53]. Here, we follow the base-$4$ numeral conventions as Ref. [53], to which we refer the reader for further implementation details.